answersLogoWhite

0

🎒

Probability

The probability of a certain event is a number expressing the likelihood that a specific event will occur, expressed as the ratio of the number of actual occurrences to the number of possible occurrences. In mathematics, it is a measure of how often an event will happen and is the basis of statistics.

14,643 Questions

In how many ways can 5 speakers be seated in a row on a stage?

Imagine the 5 seats on the stage being filled in some order. The first seat can be occupied by any one of the 5 speakers. With each choice of 5 speakers, there are four who can go for the next seat. Then, wiyth each of the 5*4 choices, there are 3 speakers left to go for the next.

Continuing in this way, the answer is 5*4*3*2*1 = 5! = 120 ways.

If one card is drawn from a deck of 52 cards what is the probability that it will either be a club or seven of hearts?

In a standard deck of 52 cards, there are 13 clubs and 1 seven of hearts.

The probability of drawing a club or the seven of hearts, then, is 14 in 52, or 7 in 26.

If you roll two dice what is the probability of getting an odd number?

The probability of getting an odd number when you roll one die is 1 in 2. In order to get an odd sum with two dice, one of them has to be odd and one of them has to be even. The probability of rolling an odd sum is still 1 in 2, since each die is unrelated in probability to the other.

What is the probablity of rolling the sum of 12 on 2 dice?

There are 36 outcomes for rolling 2 dice, and there is 1 way that a 12 can occur which is 6,6. So, the probability of rolling the sum of 12 on 2 dice is 1/36.

What is the probability of rolling the sum of 12 with 2 dice?

Since you need a specific number - the number 6 - twice, the probability is (1/6)2, which is equal to 1/36.

Since you need a specific number - the number 6 - twice, the probability is (1/6)2, which is equal to 1/36.

Since you need a specific number - the number 6 - twice, the probability is (1/6)2, which is equal to 1/36.

Since you need a specific number - the number 6 - twice, the probability is (1/6)2, which is equal to 1/36.

What is the difference between z-score and t-score?

T-scores and z-scores measure the deviation from normal. The normal for T-score is 50 with standard deviation of 10. if the score on t-score is more than 50, it means that the person scored above normal (average), and vise versa. The normal for Z-score is 0. If Z-score is above 0, then it means that person scored above normal (average), and vise versa.

What is the probability of rolling an odd number and a prime number?

With a standard 6 faced die, numbered 1 through 6, there are three odd numbers, 1, 3, and 5, of which 3 and 5 are prime, and there is one even prime, 2. The probability, then, of rolling an odd number which is prime is 1 in 3, i.e. 3 or 5 out of a possible result set of 1, 2, 3, 4, 5, and 6.

If you flip a coin 4 times how many possible outcomes are there?

There are 24 = 16 ordered outcomes, that is outcomes in which the order of the results is relevant. If not, there are 5 outcomes (0 heads, 1 head, 2 heads, 3 heads and 4 heads).

How many ways are there to select 10 balls from unlimited nunber of red white blue and green balls?

in those 10 balls we can have any combination...I am using here multinomial...

X=No. of red balls(0,1...10)

y=No. of green balls(0,1...10)

z=No. of white balls(0,1...10)

we need co-efficient of x^10 in (x^0+x^1+...+x^10)^3

that is (10+3-1)C3=12C3

Arrangement is also considered here

What is the probability of not tossing three heads with three fair coins?

The probability of something NOT happening is the complement of the probability of something happening. Since the probability that you DO have 3 heads is 1/8 (that is, 1/2 cubed), the complement is 1 - 1/8 = 7/8.

What is the probability of picking a black ace?

The probability of picking a black ace in one random draw from a normal pack of playing cards is 1/26.

You randomly select one card from a 52 card deck find the probability of selecting a nine or a two?

There are 4 nines and 4 twos, eight cards with the required outcome.

So the probability of a nine or two is 8/52 = 2/13 or 15.38%

What is the probability of selecting a prime number from 1 to 20?

There are eight prime numbers between 1 and 20.
2, 3, 5, 7, 11, 13, 17, 19

If you randomly choose in number then you have an 8 in 20 chance of selecting a prime.

The probability is selecting a prime number is 8/20 or 0.4

Why is it necessary to use a continuity correction when using a normal distribution to approximate a binomial distribution?

It is necessary to use a continuity correction when using a normal distribution to approximate a binomial distribution because the normal distribution contains real observations, while the binomial distribution contains integer observations.

How do you find out a quiz score?

You find out a quiz score by waiting for the quiz to be graded and returned to you. If it is not returned, you ask the teacher or proctor of the quiz.

If there are 50 raffle tickets in a jar and only 2 will be drawn out and you own 2 of them what chance do you have of owning both winning tickets?

2/50 x 2/50= 4/2500= 1/625

-------------------------

If the first ticket won't be replaced after it has been drawn, then your chances are 2/50*1/49=2/2450=1/1225.

What are some examples of decimals in everyday life?

hey dude somewhat right answer but read

like when you buy stuff the prices are usually decimals..

. like $3.99

your body temperature at a normal temp is 98.6 which is a decimal...